OpenStudy (anonymous):

How do I subtract 3√3i - 3i without using a calculator?

OpenStudy (anonymous):

is 'i' the imaginary number or a variable? @abcdx6

OpenStudy (anonymous):

it's an imaginary number

OpenStudy (anonymous):

I'm trying to learn how to add and subtract these kinds of problems even without the imaginary number..

OpenStudy (anonymous):

You don't need to. Just factor out the 3i from both terms: 3√3i - 3i = (√3 - 1)3i = (3√3 - 3)i. That's the final complex number. If you want to write it in the form a + bi then you can re-write that as 0 + (3√3 - 3)i where a = 0 and b = (3√3 - 3). @abcdx6

OpenStudy (anonymous):

omg thank you! :)